OpenStudy (anonymous):

Find the eleventh term of the sequence 5,9,13,17... a. 44 b. 49 c. 45 d. 53

OpenStudy (anonymous):

the asnwer is A) the go up by 4

OpenStudy (anonymous):

answer**

OpenStudy (anonymous):

oops I am sooooo sorry the answer is C) 45.

OpenStudy (cggurumanjunath):

5+4=9 9+4=13 13+4=17 17+4=21 21+4=25 25+4=29 29+4=33 33+4=37 37+4=41 41+4=45

OpenStudy (tkhunny):

First, and foremost, there is no such thing as "the eleventh term". There are infintely many answers to that question. Seriously, it can be ANYTHING you like. Second, If the question had said this, "Find AN eleventh term of the sequence 5,9,13,17... and justify your answer," that would be a much more honest question. However, there are still infinitely many responses. None more valid than any other. Third, if the question had said this, ""Find the eleventh term of the ARITHMETIC sequence 5,9,13,17...," that would be an appropriate question with a unique answer. If we number the terms, n = 1, 2, 3, 4..., we get \(Term_{n} = 5 + (n-1)\cdot 4\) \(Term_{1} = 5 + (1-1)\cdot 4 = 5 + 0\cdot 4 = 5+0 = 5\) \(Term_{2} = 5 + (2-1)\cdot 4 = 5 + 1\cdot 4 = 5+4 = 9\) \(Term_{3} = 5 + (3-1)\cdot 4 = 5 + 2\cdot 4 = 5+8 = 13\) See how that works? Find \(Term_{11}\)

OpenStudy (anonymous):

oh ok i isee :) thank you!